Given that x has a Poisson distribution with u=14, what is the probability that x=9?
Answer by jim_thompson5910(35256)

P(X = 9) = (e^(-mu)*mu^x)/(x!)

= (e^(-14)*(14)^9)/(9!)

= (0.00000083152871*20661046784)/(362880)

= 0.04734417322407


The approximate answer is 0.04734